\begin{thebibliography}{10}

\bibitem{mmseg2020}
MMSegmentation Contributors.
\newblock {MMSegmentation}: Openmmlab semantic segmentation toolbox and
  benchmark.
\newblock \url{https://github.com/open-mmlab/mmsegmentation}, 2020.

\bibitem{openseg2019}
OpenSeg Contributors.
\newblock The official pytorch implementation of ocnet series and segfix.
\newblock \url{https://github.com/openseg-group/openseg.pytorch}, 2019.

\bibitem{pytorchencoding2017}
PyTorch-Encoding Contributors.
\newblock An optimized pytorch package with cuda backend.
\newblock \url{https://github.com/zhanghang1989/PyTorch-Encoding}, 2017.

\bibitem{paszke2017automatic}
Adam Paszke, Sam Gross, Soumith Chintala, Gregory Chanan, Edward Yang, Zachary
  DeVito, Zeming Lin, Alban Desmaison, Luca Antiga, and Adam Lerer.
\newblock Automatic differentiation in pytorch.
\newblock 2017.

\bibitem{ronneberger2015u}
Olaf Ronneberger, Philipp Fischer, and Thomas Brox.
\newblock U-net: Convolutional networks for biomedical image segmentation.
\newblock In {\em Medical Image Computing and Computer-Assisted
  Intervention--MICCAI 2015: 18th International Conference, Munich, Germany,
  October 5-9, 2015, Proceedings, Part III 18}, pages 234--241. Springer, 2015.

\bibitem{bao2021beit}
Hangbo Bao, Li~Dong, Songhao Piao, and Furu Wei.
\newblock Beit: Bert pre-training of image transformers.
\newblock {\em arXiv preprint arXiv:2106.08254}, 2021.

\bibitem{devlin2018bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
\newblock Bert: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock {\em arXiv preprint arXiv:1810.04805}, 2018.

\bibitem{chu2021twins}
Xiangxiang Chu, Zhi Tian, Yuqing Wang, Bo~Zhang, Haibing Ren, Xiaolin Wei,
  Huaxia Xia, and Chunhua Shen.
\newblock Twins: Revisiting the design of spatial attention in vision
  transformers.
\newblock {\em Advances in Neural Information Processing Systems},
  34:9355--9366, 2021.

\bibitem{wu2020cgnet}
Tianyi Wu, Sheng Tang, Rui Zhang, Juan Cao, and Yongdong Zhang.
\newblock Cgnet: A light-weight context guided network for semantic
  segmentation.
\newblock {\em IEEE Transactions on Image Processing}, 30:1169--1179, 2020.

\bibitem{wang2020deep}
Jingdong Wang, Ke~Sun, Tianheng Cheng, Borui Jiang, Chaorui Deng, Yang Zhao,
  Dong Liu, Yadong Mu, Mingkui Tan, Xinggang Wang, et~al.
\newblock Deep high-resolution representation learning for visual recognition.
\newblock {\em IEEE transactions on pattern analysis and machine intelligence},
  43(10):3349--3364, 2020.

\bibitem{romera2017erfnet}
Eduardo Romera, Jos{\'e}~M Alvarez, Luis~M Bergasa, and Roberto Arroyo.
\newblock Erfnet: Efficient residual factorized convnet for real-time semantic
  segmentation.
\newblock {\em IEEE Transactions on Intelligent Transportation Systems},
  19(1):263--272, 2017.

\bibitem{he2016deep}
Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun.
\newblock Deep residual learning for image recognition.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 770--778, 2016.

\bibitem{zhang2022resnest}
Hang Zhang, Chongruo Wu, Zhongyue Zhang, Yi~Zhu, Haibin Lin, Zhi Zhang, Yue
  Sun, Tong He, Jonas Mueller, R~Manmatha, et~al.
\newblock Resnest: Split-attention networks.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 2736--2746, 2022.

\bibitem{liu2022convnet}
Zhuang Liu, Hanzi Mao, Chao-Yuan Wu, Christoph Feichtenhofer, Trevor Darrell,
  and Saining Xie.
\newblock A convnet for the 2020s.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 11976--11986, 2022.

\bibitem{poudel2019fast}
Rudra~PK Poudel, Stephan Liwicki, and Roberto Cipolla.
\newblock Fast-scnn: Fast semantic segmentation network.
\newblock {\em arXiv preprint arXiv:1902.04502}, 2019.

\bibitem{yu2018bisenet}
Changqian Yu, Jingbo Wang, Chao Peng, Changxin Gao, Gang Yu, and Nong Sang.
\newblock Bisenet: Bilateral segmentation network for real-time semantic
  segmentation.
\newblock In {\em Proceedings of the European conference on computer vision
  (ECCV)}, pages 325--341, 2018.

\bibitem{yu2021bisenet}
Changqian Yu, Changxin Gao, Jingbo Wang, Gang Yu, Chunhua Shen, and Nong Sang.
\newblock Bisenet v2: Bilateral network with guided aggregation for real-time
  semantic segmentation.
\newblock {\em International Journal of Computer Vision}, 129:3051--3068, 2021.

\bibitem{sandler2018mobilenetv2}
Mark Sandler, Andrew Howard, Menglong Zhu, Andrey Zhmoginov, and Liang-Chieh
  Chen.
\newblock Mobilenetv2: Inverted residuals and linear bottlenecks.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 4510--4520, 2018.

\bibitem{howard2019searching}
Andrew Howard, Mark Sandler, Grace Chu, Liang-Chieh Chen, Bo~Chen, Mingxing
  Tan, Weijun Wang, Yukun Zhu, Ruoming Pang, Vijay Vasudevan, et~al.
\newblock Searching for mobilenetv3.
\newblock In {\em Proceedings of the IEEE/CVF international conference on
  computer vision}, pages 1314--1324, 2019.

\bibitem{liu2021swin}
Ze~Liu, Yutong Lin, Yue Cao, Han Hu, Yixuan Wei, Zheng Zhang, Stephen Lin, and
  Baining Guo.
\newblock Swin transformer: Hierarchical vision transformer using shifted
  windows.
\newblock In {\em Proceedings of the IEEE/CVF international conference on
  computer vision}, pages 10012--10022, 2021.

\bibitem{dosovitskiy2020image}
Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn,
  Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg
  Heigold, Sylvain Gelly, et~al.
\newblock An image is worth 16x16 words: Transformers for image recognition at
  scale.
\newblock {\em arXiv preprint arXiv:2010.11929}, 2020.

\bibitem{long2015fully}
Jonathan Long, Evan Shelhamer, and Trevor Darrell.
\newblock Fully convolutional networks for semantic segmentation.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 3431--3440, 2015.

\bibitem{ruan2019devil}
Tao Ruan, Ting Liu, Zilong Huang, Yunchao Wei, Shikui Wei, and Yao Zhao.
\newblock Devil in the details: Towards accurate single and multiple human
  parsing.
\newblock In {\em Proceedings of the AAAI conference on artificial
  intelligence}, volume~33, pages 4814--4821, 2019.

\bibitem{zheng2021rethinking}
Sixiao Zheng, Jiachen Lu, Hengshuang Zhao, Xiatian Zhu, Zekun Luo, Yabiao Wang,
  Yanwei Fu, Jianfeng Feng, Tao Xiang, Philip~HS Torr, et~al.
\newblock Rethinking semantic segmentation from a sequence-to-sequence
  perspective with transformers.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 6881--6890, 2021.

\bibitem{jin2021isnet}
Zhenchao Jin, Bin Liu, Qi~Chu, and Nenghai Yu.
\newblock Isnet: Integrate image-level and semantic-level context for semantic
  segmentation.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 7189--7198, 2021.

\bibitem{zhao2018icnet}
Hengshuang Zhao, Xiaojuan Qi, Xiaoyong Shen, Jianping Shi, and Jiaya Jia.
\newblock Icnet for real-time semantic segmentation on high-resolution images.
\newblock In {\em Proceedings of the European conference on computer vision
  (ECCV)}, pages 405--420, 2018.

\bibitem{huang2019ccnet}
Zilong Huang, Xinggang Wang, Lichao Huang, Chang Huang, Yunchao Wei, and Wenyu
  Liu.
\newblock Ccnet: Criss-cross attention for semantic segmentation.
\newblock In {\em Proceedings of the IEEE/CVF international conference on
  computer vision}, pages 603--612, 2019.

\bibitem{fu2019dual}
Jun Fu, Jing Liu, Haijie Tian, Yong Li, Yongjun Bao, Zhiwei Fang, and Hanqing
  Lu.
\newblock Dual attention network for scene segmentation.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 3146--3154, 2019.

\bibitem{he2019dynamic}
Junjun He, Zhongying Deng, and Yu~Qiao.
\newblock Dynamic multi-scale filters for semantic segmentation.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 3562--3572, 2019.

\bibitem{cao2019gcnet}
Yue Cao, Jiarui Xu, Stephen Lin, Fangyun Wei, and Han Hu.
\newblock Gcnet: Non-local networks meet squeeze-excitation networks and
  beyond.
\newblock In {\em Proceedings of the IEEE/CVF international conference on
  computer vision workshops}, pages 0--0, 2019.

\bibitem{wang2018non}
Xiaolong Wang, Ross Girshick, Abhinav Gupta, and Kaiming He.
\newblock Non-local neural networks.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 7794--7803, 2018.

\bibitem{hu2018squeeze}
Jie Hu, Li~Shen, and Gang Sun.
\newblock Squeeze-and-excitation networks.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 7132--7141, 2018.

\bibitem{huang2019interlaced}
Lang Huang, Yuhui Yuan, Jianyuan Guo, Chao Zhang, Xilin Chen, and Jingdong
  Wang.
\newblock Interlaced sparse self-attention for semantic segmentation.
\newblock {\em arXiv preprint arXiv:1907.12273}, 2019.

\bibitem{zhang2018context}
Hang Zhang, Kristin Dana, Jianping Shi, Zhongyue Zhang, Xiaogang Wang, Ambrish
  Tyagi, and Amit Agrawal.
\newblock Context encoding for semantic segmentation.
\newblock In {\em Proceedings of the IEEE conference on Computer Vision and
  Pattern Recognition}, pages 7151--7160, 2018.

\bibitem{yuan2020object}
Yuhui Yuan, Xilin Chen, and Jingdong Wang.
\newblock Object-contextual representations for semantic segmentation.
\newblock In {\em Computer Vision--ECCV 2020: 16th European Conference,
  Glasgow, UK, August 23--28, 2020, Proceedings, Part VI 16}, pages 173--190.
  Springer, 2020.

\bibitem{yin2020disentangled}
Minghao Yin, Zhuliang Yao, Yue Cao, Xiu Li, Zheng Zhang, Stephen Lin, and Han
  Hu.
\newblock Disentangled non-local neural networks.
\newblock In {\em Computer Vision--ECCV 2020: 16th European Conference,
  Glasgow, UK, August 23--28, 2020, Proceedings, Part XV 16}, pages 191--207.
  Springer, 2020.

\bibitem{zhu2019asymmetric}
Zhen Zhu, Mengde Xu, Song Bai, Tengteng Huang, and Xiang Bai.
\newblock Asymmetric non-local neural networks for semantic segmentation.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 593--602, 2019.

\bibitem{li2019expectation}
Xia Li, Zhisheng Zhong, Jianlong Wu, Yibo Yang, Zhouchen Lin, and Hong Liu.
\newblock Expectation-maximization attention networks for semantic
  segmentation.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 9167--9176, 2019.

\bibitem{zhao2017pyramid}
Hengshuang Zhao, Jianping Shi, Xiaojuan Qi, Xiaogang Wang, and Jiaya Jia.
\newblock Pyramid scene parsing network.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 2881--2890, 2017.

\bibitem{zhao2018psanet}
Hengshuang Zhao, Yi~Zhang, Shu Liu, Jianping Shi, Chen~Change Loy, Dahua Lin,
  and Jiaya Jia.
\newblock Psanet: Point-wise spatial attention network for scene parsing.
\newblock In {\em Proceedings of the European conference on computer vision
  (ECCV)}, pages 267--283, 2018.

\bibitem{he2019adaptive}
Junjun He, Zhongying Deng, Lei Zhou, Yali Wang, and Yu~Qiao.
\newblock Adaptive pyramid context network for semantic segmentation.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 7519--7528, 2019.

\bibitem{wu2019fastfcn}
Huikai Wu, Junge Zhang, Kaiqi Huang, Kongming Liang, and Yizhou Yu.
\newblock Fastfcn: Rethinking dilated convolution in the backbone for semantic
  segmentation.
\newblock {\em arXiv preprint arXiv:1903.11816}, 2019.

\bibitem{xiao2018unified}
Tete Xiao, Yingcheng Liu, Bolei Zhou, Yuning Jiang, and Jian Sun.
\newblock Unified perceptual parsing for scene understanding.
\newblock In {\em Proceedings of the European conference on computer vision
  (ECCV)}, pages 418--434, 2018.

\bibitem{kirillov2020pointrend}
Alexander Kirillov, Yuxin Wu, Kaiming He, and Ross Girshick.
\newblock Pointrend: Image segmentation as rendering.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 9799--9808, 2020.

\bibitem{chen2017rethinking}
Liang-Chieh Chen, George Papandreou, Florian Schroff, and Hartwig Adam.
\newblock Rethinking atrous convolution for semantic image segmentation.
\newblock {\em arXiv preprint arXiv:1706.05587}, 2017.

\bibitem{xie2021segformer}
Enze Xie, Wenhai Wang, Zhiding Yu, Anima Anandkumar, Jose~M Alvarez, and Ping
  Luo.
\newblock Segformer: Simple and efficient design for semantic segmentation with
  transformers.
\newblock {\em Advances in Neural Information Processing Systems},
  34:12077--12090, 2021.

\bibitem{cheng2021per}
Bowen Cheng, Alex Schwing, and Alexander Kirillov.
\newblock Per-pixel classification is not all you need for semantic
  segmentation.
\newblock {\em Advances in Neural Information Processing Systems},
  34:17864--17875, 2021.

\bibitem{kirillov2019panoptic}
Alexander Kirillov, Ross Girshick, Kaiming He, and Piotr Doll{\'a}r.
\newblock Panoptic feature pyramid networks.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 6399--6408, 2019.

\bibitem{chen2018encoder}
Liang-Chieh Chen, Yukun Zhu, George Papandreou, Florian Schroff, and Hartwig
  Adam.
\newblock Encoder-decoder with atrous separable convolution for semantic image
  segmentation.
\newblock In {\em Proceedings of the European conference on computer vision
  (ECCV)}, pages 801--818, 2018.

\bibitem{jin2021mining}
Zhenchao Jin, Tao Gong, Dongdong Yu, Qi~Chu, Jian Wang, Changhu Wang, and Jie
  Shao.
\newblock Mining contextual information beyond image for semantic segmentation.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 7231--7241, 2021.

\bibitem{jin2022mcibi++}
Zhenchao Jin, Dongdong Yu, Zehuan Yuan, and Lequan Yu.
\newblock Mcibi++: Soft mining contextual information beyond image for semantic
  segmentation.
\newblock {\em IEEE Transactions on Pattern Analysis and Machine Intelligence},
  2022.

\bibitem{micikevicius2017mixed}
Paulius Micikevicius, Sharan Narang, Jonah Alben, Gregory Diamos, Erich Elsen,
  David Garcia, Boris Ginsburg, Michael Houston, Oleksii Kuchaiev, Ganesh
  Venkatesh, et~al.
\newblock Mixed precision training.
\newblock {\em arXiv preprint arXiv:1710.03740}, 2017.

\bibitem{gong2017look}
Ke~Gong, Xiaodan Liang, Dongyu Zhang, Xiaohui Shen, and Liang Lin.
\newblock Look into person: Self-supervised structure-sensitive learning and a
  new benchmark for human parsing.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 932--940, 2017.

\bibitem{zhou2017scene}
Bolei Zhou, Hang Zhao, Xavier Puig, Sanja Fidler, Adela Barriuso, and Antonio
  Torralba.
\newblock Scene parsing through ade20k dataset.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 633--641, 2017.

\bibitem{cordts2016cityscapes}
Marius Cordts, Mohamed Omran, Sebastian Ramos, Timo Rehfeld, Markus Enzweiler,
  Rodrigo Benenson, Uwe Franke, Stefan Roth, and Bernt Schiele.
\newblock The cityscapes dataset for semantic urban scene understanding.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 3213--3223, 2016.

\bibitem{caesar2018coco}
Holger Caesar, Jasper Uijlings, and Vittorio Ferrari.
\newblock Coco-stuff: Thing and stuff classes in context.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 1209--1218, 2018.

\bibitem{mottaghi2014role}
Roozbeh Mottaghi, Xianjie Chen, Xiaobai Liu, Nam-Gyu Cho, Seong-Whan Lee, Sanja
  Fidler, Raquel Urtasun, and Alan Yuille.
\newblock The role of context for object detection and semantic segmentation in
  the wild.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 891--898, 2014.

\bibitem{miao2021vspw}
Jiaxu Miao, Yunchao Wei, Yu~Wu, Chen Liang, Guangrui Li, and Yi~Yang.
\newblock Vspw: A large-scale dataset for video scene parsing in the wild.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 4133--4143, 2021.

\end{thebibliography}

\documentclass{article}

\usepackage{PRIMEarxiv}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{lipsum}
\usepackage{fancyhdr}       
\usepackage{graphicx}       
\usepackage{float}
\graphicspath{{media/}}     

\title{SSSegmenation: An Open Source Supervised Semantic Segmentation Toolbox Based on PyTorch}

\author{
  Zhenchao Jin \\
  The University of Hong Kong \\
  \texttt{blwx96@connect.hku.hk} 
}

\begin{document}
\maketitle

\begin{abstract}
  This paper presents SSSegmenation, which is an open source supervised semantic image segmentation toolbox based on PyTorch.
  The design of this toolbox is motivated by MMSegmentation while it is easier to use because of fewer dependencies and achieves superior segmentation performance under a comparable training and testing setup.
  Moreover, the toolbox also provides plenty of trained weights for popular and contemporary semantic segmentation methods, including Deeplab, PSPNet, OCRNet, MaskFormer, \emph{etc}.
  We expect that this toolbox can contribute to the future development of semantic segmentation.
  Codes and model zoos are available at \href{https://github.com/SegmentationBLWX/sssegmentation/}{SSSegmenation}.
\end{abstract}


\section{Introduction}

Semantic image segmentation is a task of grouping pixels with different semantic categories.
Prior to SSSegmenation, many efforts have been made to provide high-quality codebases and unified benchmarks for semantic segmentation, \emph{e.g.}, MMSegmentation \cite{mmseg2020}, OpenSeg \cite{openseg2019} and PyTorch-Encoding \cite{pytorchencoding2017}.
Despite this, most of the codebases have not been updated for a long time and thus, lack support for the state-of-the-art segmentation algorithms.
Furthermore, the actively updated MMSegmentation toolbox contains complex mmlab-series dependencies, which is sometimes unfriendly to the researchers.
To reduce the learning cost of using a segmentation toolbox and provide high-quality re-implementation of the state-of-the-art segmentation approaches, we build SSSegmenation, an open source supervised semantic image segmentation toolbox based on PyTorch \cite{paszke2017automatic}.
The major features of SSSegmenation can be summarized as follows,
\begin{itemize}
  \item \textbf{High Performance.} The performance of re-implemented semantic segmentation algorithms is better than or comparable to other codebases.
  \item \textbf{Unified Benchmark.} Like conventional MMSegmentation \cite{mmseg2020}, various segmentation methods are unified into several specific modules. Benefiting from this design, SSSegmentation can integrate a great deal of popular and contemporary semantic segmentation frameworks and then, train and test them on unified benchmarks.
  \item \textbf{Fewer Dependencies.} SSSegmenation tries its best to avoid introducing more dependencies when reproducing novel semantic segmentation approaches.
\end{itemize}
SSSegmenation project is under active development and, we expect that this toolbox can facilitate the future development of semantic image segmentation.

\section{Supported Frameworks}

SSSegmenation supports abounding backbone networks (\emph{i.e.}, encoder) and segmentors (\emph{i.e.}, decoder).
The integrated backbone networks and segmentors in SSSegmenation are listed in this section.
Note that, most of the re-implementation algorithms in SSSegmenation are modified from the released codes in official repositories or MMSegmentation \cite{mmseg2020}. 

\subsection{Supported Backbone Networks}

\begin{itemize}
  \item \textbf{UNet \cite{ronneberger2015u}.} A classic backbone network in medical image segmentation, which is published in MICCAI 2015.
  \item \textbf{BEiT \cite{bao2021beit}.} A self-supervised bidirectional encoder vision representation model inspired by BERT \cite{devlin2018bert}, and it is accepted by ICLR 2022.
  \item \textbf{Twins \cite{chu2021twins}.} A variant of vision transformer architectures for visual tasks, which appears in NeurIPS 2021.
  \item \textbf{CGNet \cite{wu2020cgnet}.} A lightweight and efficient network for semantic segmentation, which is accepted by TIP 2020.
  \item \textbf{HRNet \cite{wang2020deep}.} A high-resolution representation learning network, and it appears in TPAMI 2020.
  \item \textbf{ERFNet \cite{romera2017erfnet}.} An efficient residual factorized convolution network for real-time semantic segmentation, which is published in T-ITS 2017.
  \item \textbf{ResNet \cite{he2016deep}.} A classic backbone network in image classification, which is accepted by CVPR 2016.
  \item \textbf{ResNeSt \cite{zhang2022resnest}.} A split-attention network for visual tasks, which appears in CVPRW 2022.
  \item \textbf{ConvNeXt \cite{liu2022convnet}.} A ViT-inspired convolution network, and it is published in CVPR 2022.
  \item \textbf{FastSCNN \cite{poudel2019fast}.} A fast segmentation convolutional neural network reported in	ArXiv 2019.
  \item \textbf{BiSeNetV1 \cite{yu2018bisenet}.} A real-time bilateral segmentation network, which is accepted by ECCV 2018.
  \item \textbf{BiSeNetV2 \cite{yu2021bisenet}.} A bilateral network with guided aggregation for real-time semantic segmentation, which is published in IJCV 2021.
  \item \textbf{MobileNetV2 \cite{sandler2018mobilenetv2}.} A widely used lightweight and real-time convolution network for various visual applications, which is accepted by CVPR 2018.
  \item \textbf{MobileNetV3 \cite{howard2019searching}.} A lightweight and real-time convolution network based on a combination of complementary search techniques as well as a novel architecture design, which appears in ICCV 2019.
  \item \textbf{SwinTransformer \cite{liu2021swin}.} A widely used hierarchical vision transformer architecture using shifted windows for visual tasks, which is published in ICCV 2021.
  \item \textbf{VisionTransformer \cite{dosovitskiy2020image}.} The first vision transformer architecture, which is accepted by IClR 2021.
\end{itemize}

\subsection{Supported Segmentors}

\begin{itemize}
  \item \textbf{FCN \cite{long2015fully}.} A conventional encoder-decoder segmentation framework, and it is published in TPAMI 2017.
  \item \textbf{CE2P \cite{ruan2019devil}.} A classic human parsing framework, which appears in AAAI 2019.
  \item \textbf{SETR \cite{zheng2021rethinking}.} A pure transformer structure (\emph{i.e.}, without convolution and resolution reduction) to encode an image as a sequence of patches, which is proposed in CVPR 2021.
  \item \textbf{ISNet \cite{jin2021isnet}.} A high-performance semantic segmentation framework utilizing image-level and semantic-level contextual information to augment pixel representations, which is proposed in ICCV 2021.
  \item \textbf{ICNet \cite{zhao2018icnet}.} An image cascade network that incorporates multi-resolution branches under proper label guidance to reduce a large portion of computation for pixel-wise label inference, which is introduced in ECCV 2018.
  \item \textbf{CCNet \cite{huang2019ccnet}.} A criss-cross network for effectively and efficiently obtaining full-image contextual information, which is accepted by ICCV 2019.
  \item \textbf{DANet \cite{fu2019dual}.} A dual attention network to adaptively integrate local features with their global dependencies, which is published in	CVPR 2019.
  \item \textbf{DMNet \cite{he2019dynamic}.} A dynamic multi-scale neural network to adaptively aggregate multi-scale contents for predicting pixel-level semantic labels, which is proposed in ICCV 2019.
  \item \textbf{GCNet \cite{cao2019gcnet}.} A segmentation framework that unifies non-local block \cite{wang2018non} and squeeze-excitation network \cite{hu2018squeeze}  into a three-step general framework for global context modeling, which is proposed in TPAMI 2020.
  \item \textbf{ISANet \cite{huang2019interlaced}.} An interlaced sparse self-attention approach to improve the efficiency of self-attention mechanism for semantic segmentation, which is accepted by IJCV 2021.
  \item \textbf{EncNet \cite{zhang2018context}.} A context encoding module to explore the impact of global contextual information in semantic segmentation, which is presented in CVPR 2018.
  \item \textbf{OCRNet \cite{yuan2020object}.} A semantic segmentation framework that emphasizes utilizing object-contextual representations to enhance pixel representations, which is proposed in ECCV 2020.
  \item \textbf{DNLNet \cite{yin2020disentangled}.} A disentangled non-local block to facilitate learning for both the whitened pairwise term and the unary term in semantic segmentation, which appears in ECCV 2020.
  \item \textbf{ANNNet \cite{zhu2019asymmetric}.} An asymmetric non-local neural network (consisting of asymmetric pyramid non-local block and asymmetric fusion non-local block) for semantic segmentation, which is accepted by ICCV 2019.
  \item \textbf{EMANet \cite{li2019expectation}.} An expectation-maximization attention module for aggregating contextual information in a robust way, which is proposed in ICCV 2019.
  \item \textbf{PSPNet \cite{zhao2017pyramid}.} A pyramid pooling module is designed to capture multi-scale contextual information, which is accepted by CVPR 2017.
  \item \textbf{PSANet \cite{zhao2018psanet}.} A point-wise spatial attention network to relax the local neighborhood constraint for semantic segmentation, which is introduced in CVPR 2017.
  \item \textbf{APCNet \cite{he2019adaptive}.} An adaptive pyramid context network for semantic segmentation, accepted by CVPR 2019.
  \item \textbf{FastFCN \cite{wu2019fastfcn}.} A joint pyramid upsampling module to formulate the task of extracting high-resolution feature maps into a joint upsampling problem, which is reported in	ArXiv 2019.
  \item \textbf{UPerNet \cite{xiao2018unified}.} A hierarchical structure to learn from heterogeneous data from multiple image datasets in semantic segmentation, which is proposed in ECCV 2018.
  \item \textbf{PointRend \cite{kirillov2020pointrend}.}  Taking image segmentation as a rendering problem, proposed in CVPR 2020.
  \item \textbf{Deeplabv3 \cite{chen2017rethinking}.} A segmentation framework that uses atrous convolution in cascade or in parallel to capture multi-scale context by adopting multiple atrous rates, which is introduced in CVPR 2017.
  \item \textbf{Segformer \cite{xie2021segformer}.} A simple, efficient yet powerful semantic segmentation framework which unifies transformers with lightweight multilayer perceptron (MLP) decoders, proposed in NeurIPS 2021.
  \item \textbf{MaskFormer \cite{cheng2021per}.} A simple mask classification model that formulates semantic segmentation as a per-pixel classification task, which is published in NeurIPS 2021.
  \item \textbf{SemanticFPN \cite{kirillov2019panoptic}.} A panoptic feature pyramid network applied in semantic segmentation, which is proposed in CVPR 2019.
  \item \textbf{NonLocalNet \cite{wang2018non}.} A non-local block to capture long-range dependencies for semantic segmentation, which is accepted by CVPR 2018.
  \item \textbf{Deeplabv3Plus \cite{chen2018encoder}.} An extended version of Deeplabv3, which is published in ECCV 2018.
  \item \textbf{MCIBI \cite{jin2021mining}.} A semantic segmentation framework targeting mining contextual information beyond image, which is accepted by ICCV 2021.
  \item \textbf{MCIBI++ \cite{jin2022mcibi++}.} An extended version of MCIBI, which is published in TPAMI 2022.
  \item \textbf{Mixed Precision (FP16) Training \cite{micikevicius2017mixed}.} Mixed precision training aims at training deep neural networks using half precision floating point (FP16) numbers, and it is proposed in ArXiv 2017.
\end{itemize}

\section{Benchmarks}

The segmentation frameworks integrated into SSSegmenation can be trained on various popular benchmark datasets, including LIP \cite{gong2017look}, ADE20k \cite{zhou2017scene}, CityScapes \cite{cordts2016cityscapes}, COCO Stuff \cite{caesar2018coco}, Pascal Context \cite{mottaghi2014role}, VSPW \cite{miao2021vspw}, to name a few.
Table \ref{tab:reportedperformance} shows the performance of various segmentation methods on Pascal VOC, ADE20K and Cityscapes datasets.
The readers can refer to the official repository of SSSegmenation for training details and more segmentation results.

\begin{table}[t]
\centering
\caption{
  The performance of different segmentation approaches on three popular benchmark datasets.
  All results are obtained under a single-scale setting without extra tricks, \emph{e.g.}, flipping and multi-scale testing.
}\label{tab:reportedperformance}
\resizebox{1.00\textwidth}{!}{
    \begin{tabular}{c|c|c|c|c|c}
    \toprule
    Method                               &Backbone                     &Stride        &ADE20K (\emph{train / val})           &Pascal VOC (\emph{train / test})     &Cityscapes (\emph{train / val})      \\
    \hline       
    ANNNet                               &ResNet-50                    &$8\times$     &41.75                                 &76.68                                &78.36                                \\
    ANNNet                               &ResNet-101                   &$8\times$     &43.98                                 &78.15                                &79.34                                \\
    \hline
    APCNet                               &ResNet-50                    &$8\times$     &43.47                                 &76.97                                &79.02                                \\
    APCNet                               &ResNet-101                   &$8\times$     &45.74                                 &78.99                                &79.71                                \\
    \hline
    CCNet                                &ResNet-50                    &$8\times$     &42.47                                 &77.43                                &79.15                                \\
    CCNet                                &ResNet-101                   &$8\times$     &44.00                                 &78.02                                &80.08                                \\
    \hline
    DANet                                &ResNet-50                    &$8\times$     &42.90                                 &76.39                                &79.47                                \\
    DANet                                &ResNet-101                   &$8\times$     &44.37                                 &77.97                                &80.55                                \\
    \hline
    Deeplabv3                            &ResNet-50                    &$8\times$     &43.19                                 &77.72                                &79.62                                \\
    Deeplabv3                            &ResNet-101                   &$8\times$     &45.16                                 &79.52                                &80.28                                \\
    \hline
    Deeplabv3Plus                        &ResNet-50                    &$8\times$     &44.51                                 &77.43                                &80.38                                \\
    Deeplabv3Plus                        &ResNet-101                   &$8\times$     &45.72                                 &79.19                                &81.09                                \\
    Deeplabv3Plus                        &ResNeSt-101                  &$8\times$     &46.48                                 &79.76                                &80.30                                \\
    \hline
    DMNet                                &ResNet-50                    &$8\times$     &43.54                                 &77.38                                &79.17                                \\
    DMNet                                &ResNet-101                   &$8\times$     &45.53                                 &79.15                                &79.90                                \\
    \hline
    DNLNet                               &ResNet-50                    &$8\times$     &43.50                                 &76.73                                &79.75                                \\
    DNLNet                               &ResNet-101                   &$8\times$     &44.88                                 &78.37                                &80.64                                \\
    \hline
    EMANet                               &ResNet-50                    &$8\times$     &41.77                                 &75.29                                &77.96                                \\
    EMANet                               &ResNet-101                   &$8\times$     &44.39                                 &76.43                                &79.54                                \\
    \hline
    EncNet                               &ResNet-50                    &$8\times$     &40.60                                 &75.53                                &77.98                                \\
    EncNet                               &ResNet-101                   &$8\times$     &43.43                                 &77.61                                &78.70                                \\
    \hline
    GCNet                                &ResNet-50                    &$8\times$     &42.53                                 &76.50                                &78.69                                \\
    GCNet                                &ResNet-101                   &$8\times$     &44.19                                 &78.81                                &79.93                                \\
    \hline
    ICNet                                &ResNet-50                    &$8\times$     &-                                     &-                                    &76.60                                \\
    ICNet                                &ResNet-101                   &$8\times$     &-                                     &-                                    &76.27                                \\
    \hline
    ISANet                               &ResNet-50                    &$8\times$     &42.60                                 &76.99                                &79.34                                \\
    ISANet                               &ResNet-101                   &$8\times$     &44.08                                 &78.60                                &80.58                                \\
    \hline
    ISNet                                &ResNet-50                    &$8\times$     &44.22                                 &-                                    &79.32                                \\
    ISNet                                &ResNet-101                   &$8\times$     &45.92                                 &-                                    &80.56                                \\
    \hline
    Deeplabv3+MCIBI                      &ResNet-50                    &$8\times$     &44.39                                 &-                                    &-                                    \\
    Deeplabv3+MCIBI                      &ResNet-101                   &$8\times$     &45.66                                 &-                                    &-                                    \\
    \hline
    UPerNet+MCIBI++                      &ResNet-50                    &$8\times$     &44.30                                 &79.48	                               &-                                    \\
    UPerNet+MCIBI++                      &ResNet-101                   &$8\times$     &46.39                                 &80.42                                &-                                    \\
    \hline
    NonLocalNet                          &ResNet-50                    &$8\times$     &42.15                                 &77.08                                &78.34                                \\
    NonLocalNet                          &ResNet-101                   &$8\times$     &44.49                                 &78.89                                &80.42                                \\
    \hline
    OCRNet                               &ResNet-50                    &$8\times$     &42.47                                 &76.75                                &79.40                                \\
    OCRNet                               &ResNet-101                   &$8\times$     &43.99                                 &78.82                                &80.61                                \\
    OCRNet                               &HRNetV2p-W48                 &$4\times$     &44.03                                 &77.60                                &81.44                                \\
    \hline
    PointRend                            &ResNet-50                    &$8\times$     &37.80                                 &69.84                                &76.89                                \\
    PointRend                            &ResNet-101                   &$8\times$     &40.26                                 &72.31                                &78.80                                \\
    \hline
    PSANet                               &ResNet-50                    &$8\times$     &41.99                                 &77.06                                &78.88                                \\
    PSANet                               &ResNet-101                   &$8\times$     &43.85                                 &78.97                                &79.65                                \\
    \hline
    PSPNet                               &ResNet-50                    &$8\times$     &42.64                                 &77.93                                &79.05                                \\
    PSPNet                               &ResNet-101                   &$8\times$     &44.55                                 &79.04                                &79.94                                \\
    \hline
    FCN                                  &ResNet-50                    &$8\times$     &36.96                                 &67.80                                &75.16                                \\
    FCN                                  &ResNet-101                   &$8\times$     &41.22                                 &70.59                                &76.31                                \\
    FCN                                  &BiSeNetV1+ResNet-18          &$32\times$    &-                                     &-                                    &75.76                                \\
    FCN                                  &BiSeNetV1+ResNet-50          &$32\times$    &-                                     &-                                    &77.78                                \\
    FCN                                  &BiSeNetV2                    &$32\times$    &-                                     &-                                    &74.62                                \\
    FCN                                  &CGNet-M3N21                  &$8\times$     &-                                     &-                                    &68.53                                \\
    FCN                                  &ERFNet                       &$8\times$     &-                                     &-                                    &76.44                                \\
    \hline
    UPerNet                              &ResNet-50                    &$8\times$     &43.02                                 &76.86                                &79.08                                \\
    UPerNet                              &ResNet-101                   &$8\times$     &44.92                                 &79.13                                &80.39                                \\
    UPerNet                              &BEiT-B                       &$16\times$    &53.22                                 &-                                    &-                                    \\
    UPerNet                              &BEiT-L                       &$16\times$    &56.52                                 &-                                    &-                                    \\
    UPerNet                              &ConvNeXt-L-21k               &$8\times$     &53.41                                 &-                                    &-                                    \\
    UPerNet                              &ConvNeXt-XL-21k              &$8\times$     &53.68                                 &-                                    &-                                    \\
    UPerNet                              &Swin-Small                   &$32\times$    &48.39                                 &-                                    &-                                    \\
    UPerNet                              &Swin-Base                    &$32\times$    &51.02                                 &-                                    &-                                    \\
    UPerNet                              &SVT-Large                    &$32\times$    &49.80                                 &-                                    &-                                    \\
    UPerNet                              &PCPVT-Large                  &$32\times$    &49.35                                 &-                                    &-                                    \\
    \hline
    MaskFormer                           &Swin-Small                   &$32\times$    &49.91                                 &-                                    &-                                    \\
    MaskFormer                           &Swin-Base                    &$32\times$    &53.22                                 &-                                    &-                                    \\
    \hline
    SETR-Naive                           &ViT-Large                    &$32\times$    &48.43                                 &-                                    &-                                    \\
    SETR-PUP                             &ViT-Large                    &$32\times$    &48.51                                 &-                                    &-                                    \\
    SETR-MLA                             &ViT-Large                    &$32\times$    &49.61                                 &-                                    &-                                    \\
    \bottomrule
\end{tabular}}
\vspace{-0.4cm}
\end{table}

\section{Code Structure of SSSegmenation}

SSSegmenation mainly consists of \emph{dataset module}, \emph{parallel module}, \emph{backbone module}, \emph{segmentor module}, \emph{loss module}, \emph{optimizer module} and \emph{scheduler module}.
Specifically, \emph{dataset module} is defined to load training and testing images.
\emph{parallel module} simply call the $nn.parallel.DistributedDataParallel$ function in PyTorch to achieve distributed training.
\emph{backbone module} is employed to define the encoder structure while \emph{segmentor module} is leveraged to define the decoder structure.
\emph{loss module} is introduced to define the objective functions for semantic segmentation.
\emph{optimizer module} and \emph{scheduler module} are designed for optimizing the network parameters through back propagation algorithm.
Furthermore, it is also convenient to integrate more segmentation algorithms into SSSegmenation since we only need to define our algorithm in the corresponding module folder and import it in the $builder.py$ in the folder.

The minimum requirement for utilizing SSSegmenation is a CUDA development environment with PyTorch and some necessary but easy-to-installed packages, including \emph{chainercv}, \emph{pillow}, \emph{opencv-python}, \emph{pandas}, \emph{cython}, \emph{numpy}, \emph{scipy}, \emph{tqdm} and \emph{argparse}.
Also, you can install \emph{timm}, \emph{mmcv} and other packages to extend the functions of SSSegmenation.

In summary, SSSegmenation is a simple yet effective codebase for semantic image segmentation.
We hope this toolbox can further advance researches on semantic image segmentation.

\bibliographystyle{unsrt}  
\bibliography{references}

\end{document}